OpenStudy (anonymous):

What is the slope of a line that passes through the point (-2, 3) and is parallel to a line that passes through (3, 7) and (-2, -8)?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

yep

OpenStudy (phi):

this is a bit of a trick. The point (-2,3) is irrelevant. You must find the slope of the line through the points (3,7) and (-2,-8) change in y over change in x: \[ \frac{-8 -7}{-2-3} = \frac{-15}{-5} = 3 \]

OpenStudy (anonymous):

loll yeah sry :P i m kinda disturbed